# Equivariant Ehrhart theory, commutative algebra and invariant triangulations of polytopes

### v2

@@ 1. Introduction
Ehrhart theory is the study of the enumeration of lattice points in lattice polytopes. Equivariant Ehrhart theory is a generalization of Ehrhart theory that takes into account the action of a finite group acting via affine transformations on the underlying lattice and preserving the polytope. It was introduced in [Sta11a] and studied in [ASVM20, ASVM21, EKS23, CHK23]. The goal of this paper is to further develop equivariant Ehrhart theory in part by establishing a connection with commutative algebra that generalizes the known connection between commutative algebra and Ehrhart theory. In what may be considered either a feature or a bug, in contrast to the non-equivariant case, the connection with commutative algebra may or may not hold, and much of the paper is devoted to the question of when it holds. As we will see, this is related to the question of when a lattice polytope admits an invariant lattice triangulation.
Before stating our results, we first recall some background on equivariant Ehrhart theory. Let $M\cong\mathbb{Z}^{d}$ be a lattice of rank $d$ and write $M_{k}:=M\otimes_{\mathbb{Z}}k$ for a field $k$. Let $P\subset M_{\mathbb{R}}$ be a $d$-dimensional lattice polytope. Let $\operatorname{Aff}(M)$ be the group of affine transformations of $M$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation of $G$ (see Section 2.2 and Section 2.3)). We assume that $P$ is $G$-invariant. Let $R(G)$ be the complex representation ring of $G$ (see Section 2.3). Then $R(G)$ is the free $\mathbb{Z}$-module with basis given by the classes $[V]$ in $R(G)$, as $V$ varies over all irreducible finite-dimensional $\mathbb{C}G$-modules, and an element of $R(G)$ with nonnegative integer coefficients with respect to this basis is called effective. We say that a power series in $R(G)[[t]]$ is effective if all its coefficients are effective.
@@
Above, we let $\widetilde{M}=M\oplus\mathbb{Z}$ and consider $\operatorname{Aff}(M)$ as the subgroup of $\operatorname{GL}(\widetilde{M})$ consisting of linear transformations that preserve projection onto the last coordinate $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$. In particular, we consider $\widetilde{M}_{\mathbb{C}}$ as a $\mathbb{C}[G]$-module. If $V$ is a finite-dimensional $\mathbb{C}[G]$-module, we write $\det(I-Vt):=\sum_{i}(-1)^{i}[\bigwedge^{i}V]t^{i}\in R(G)[t]$ (see (2)). A priori, $h^{*}(P,\rho;t)$ depends on both $\operatorname{Ehr}(P,\rho;t)$ and the complex representation $\widetilde{M}_{\mathbb{C}}$ of $G$, but, in fact, $\operatorname{Ehr}(P,\rho;t)$ precisely encodes both $h^{*}(P,\rho;t)$ and $\widetilde{M}_{\mathbb{C}}$ (see Lemma 2.14).
@@
(see [Sta11a, Theorem 7.7]) Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Consider the following conditions:
@@
Then the following implications hold: (1) normal-$\Rightarrow$ (2) normal-$\Rightarrow$ (3).
@@
We now present Theorem 1.4, a generalization of Theorem 1.1 that is the main result of the paper. Let $C_{P}\subset\widetilde{M}_{\mathbb{R}}$ be the cone generated by $P\times\{1\}$, and consider the graded semigroup algebra $S_{P}:=\mathbb{C}[C_{P}\cap\widetilde{M}]$, with grading induced by $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$. For any positive integer $m$, there is a natural bijection between elements of $C_{P}\cap\operatorname{pr}^{-1}(m)$ and elements of $P\cap\frac{1}{m}M$, taking $(u,m)\mapsto\frac{1}{m}u$. In particular, $G$ acts on $S_{P}$ as a graded $\mathbb{C}$-algebra and $\operatorname{Ehr}(P,\rho;t)$ is the equivariant Hilbert series of $S_{P}$ (see Section 4.1). By a result of Hochster [Hoc72], $S_{P}$ is a Cohen-Macaulay ring. A linear system of parameters (l.s.o.p.) is a collection of $d+1$ homogeneous elements $F_{1},\ldots,F_{d+1}$ of degree $1$ in $S_{P}$ such that $S_{P}/(F_{1},\ldots,F_{d+1})$ is a finite dimensional $\mathbb{C}$-vector space (see Definition 4.3 and Theorem 4.4). If $F_{1},\ldots,F_{d+1}$ is a l.s.o.p. for $S_{P}$, then the ideal $(F_{1},\ldots,F_{d+1})$ is $G$-invariant if and only if the $\mathbb{C}$-vector space $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}$ is $G$-invariant if and only if we may consider $[\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]$ as an element of $R(G)$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $\widetilde{M}=M\oplus\mathbb{Z}$ and let $C_{P}$ be the cone over $P\times\{1\}$ in $\widetilde{M}_{\mathbb{R}}$ with corresponding graded semigroup algebra $S_{P}=\mathbb{C}[C_{P}\cap\widetilde{M}]$. Consider the following conditions:
@@
    There exists a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{P}$ such that $[\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}]$ in $R(G)$.
@@
Then the following implications hold: (1) normal-$\Rightarrow$ (2) normal-$\Rightarrow$ (3) normal-$\Rightarrow$ (4) normal-$\Rightarrow$ (5). Moreover, if there exists a $G$-invariant (not necessarily regular) lattice triangulation of $P$, then (4) holds.
When $G$ acts trivially, all the conditions of Theorem 1.4 are well-known to hold. See, for example, [DLRS10, Lemma 2.3.15] for condition (1) and [Sta76, Sta80] for conditions (3) and (4). We observe that condition (1) in Theorem 1.4 is often easier to verify in practice than conditions (2) or (3). In Proposition 3.12 we give an explicit criterion to guarantee that (1) holds i.e. a $G$-invariant regular lattice triangulation of $P$ exists. Using the notion of a translative action studied in [DR18], [DD21], [BD22] and [DD23] (see Definition 3.13) we deduce two corollaries for refining and gluing invariant lattice polyhedral subdivisions to construct invariant lattice triangulations (see Corollary 3.17 and Corollary 3.18 respectively). For example, we deduce that (1) holds when $G=\mathbb{Z}/2\mathbb{Z}$ acts on a centrally symmetric lattice polytope (see Example 3.20). On the other hand, we can view condition (5) in Theorem 1.4 as an obstruction that can be used to construct examples where conditions (1)-(3) fail.
For $G$-invariant lattice triangulations with some special additional structure, combinatorial formulas for the coefficients of $h^{*}(P,\rho;t)$ as explicit permutation representations were given in [EKS23, Theorem 3.30] (see, also, [EKS23, Theorem 3.32]). When the action of $G$ is translative and there exists a $G$-invariant unimodular lattice triangulation of $P$, D’alì and Delucchi showed in independent work that the equivariant Ehrhart series can be expressed as a rational function with denominator $(1-t)^{d+1}$ and gave a criterion which guarantees that the corresponding numerator is effective [DD23, Theorem 3.8], with applications to order polytopes and Lipschitz polytopes of posets ([DD23, Theorem 4.5] and [DD23, Theorem 5.3] respectively).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $N$ be a positive integer divisible by $|G|$. Then there exists a $G$-invariant regular triangulation of $P$ with vertices in $\frac{1}{N}M$.
@@
The invariant $h^{*}_{N}(P,\rho;t)$ will be studied in a more general context in [Sta23]. On the one hand, the definition is motivated by the non-equivariant case when $h^{*}_{N}(P,\rho;t)$ has an interpretation as the weighted Ehrhart polynomial of the $N$th dilate of the pyramid over $P$ (see Lemma 2.33, [Sta08] and [Sta23]). On the other hand, the definition is also motivated by the Ehrhart theory of rational polytopes where denominators with factors of the form $1-t^{N}$ (rather than $1-t$) naturally appear (see Remark 2.2). In fact, we leave it as an open problem to extend the theory of this paper to the case when $P$ is a rational (rather than a lattice) polytope (cf. [CHK23]). We recover the invariant $h^{*}(P,\rho_{N};t)$ from $h_{N}^{*}(P,\rho;t)$ by extracting all terms with integer (rather than just rational) exponents in $t$ (see Lemma 2.33).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $N$ be a positive integer. Consider the following conditions:
@@
Then the following implications hold: (1) normal-$\Rightarrow$ (2) normal-$\Rightarrow$ (3).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $Q\subset P$ be a $G$-invariant (not necessarily full dimensional) lattice polytope. Suppose there exists a $G$-invariant lattice triangulation of $P$ that restricts to a $G$-invariant lattice triangulation of $Q$. Then $h^{*}(Q,\rho_{Q};t)\leq h^{*}(P,\rho;t)$.
@@
Erratum to [Sta11a]: We mention three errors in [Sta11a], ordered from most to least serious. Firstly, [Sta11a, Remark 7.8], together with [Sta11a, Remark 7.8], give a criterion that guarantees that condition (2) in Theorem 1.4 holds. However, Example 3.22 shows that this is false. The fix is that in equation $(6)$ in [Sta11a, Remark 7.8]$G_{Q}$-orbit’ needs to be replaced by ‘$G$-orbit’. Alternatively, we provide Corollary 3.17 and Corollary 3.18, which together with Theorem 1.4 may act as a replacement (with stronger hypotheses). As a consequence of this error, some of the proofs of subsequent statements are incomplete: [Sta11a, Corollary 7.10] (we do not have a proof or a counterexample), [Sta11a, Corollary 7.12] (the statement is correct by Proposition 3.24), [Sta11a, Corollary 7.14] (the statement is correct by Theorem 1.4 and Theorem 1.5), [Sta11a, Example 7.16] (we do not have a proof or a counterexample). In the cases above where we do not have a counterexample, we do have reason to expect that the statements are true. Unfortunately, the proof of [EKS23, Theorem 3.52] references [Sta11a, Corollary 7.10] and hence is incomplete (the version of Corollary 7.10 stated in [EKS23, Corollary 2.10] is false by Example 3.22). However, the statement of [EKS23, Theorem 3.52] is true and Example 3.23 below provides a generalization (using Theorem 1.4).
Secondly, in the statement of condition (2) we consider a polynomial $f\in\mathbb{C}[M]$ that is $G$-invariant. This implies that the hypersurface $\{f=0\}\subset\operatorname{Spec}\mathbb{C}[M]$ is $G$-invariant. However, the converse is false. More precisely, there may be a nontrivial character $\chi:G\to\mathbb{C}$ such that $g\cdot f=\chi(g)f$ for all $g$ in $G$, and then $\{f=0\}$ is $G$-invariant but $f$ is not. In particular, references to $G$-invariant hypersurfaces in [Sta11a] and [Sta11b] should be references to $G$-invariant polynomials. This issue will be treated in more detail in [Sta23].
@@ 2.1. Ehrhart theory of rational polytopes
for some $10\leq\lambda_{k}<$ 1, $m_{i},m_{i}^{\prime}\in\mathbb{Z}$ and $\beta\in\mathbb{R}$. Note that $\lambda_{k}$ is determined by $\lambda_{k}=\beta a_{i}=-\beta a_{i}\in\mathbb{R}/\mathbb{Z}$ for any $i\in I_{k}$. Since $\gcd(a_{1},\ldots,a_{r})=1$, the condition $2\beta a_{i}=0\in\mathbb{R}/\mathbb{Z}$ implies that either $\beta=0\in\mathbb{R}/\mathbb{Z}$ or $\beta=\frac{1}{2}\in\mathbb{R}/\mathbb{Z}$. If $\beta=0\in\mathbb{R}/\mathbb{Z}$, then $\lambda_{k}=0$ for all $k$ and $\bar{u}=0$. Assume that $\beta=\frac{1}{2}\in\mathbb{R}/\mathbb{Z}$. Then $\lambda_{k}=\frac{1}{2}$ if $a_{i}$ is odd for some/any $i\in I_{k}$, and $\lambda_{k}=0$ otherwise. The only solution is $u=\frac{1}{2}\sum_{i:a_{i}\textrm{ odd }}(e_{i}+f_{i})$.

### v3

@@ 1. Introduction
The presence of a group action can lead to subtle questions. For example, given a lattice polytope $P$, there always exists a regular lattice triangulation of $P$. On the other hand, if a group $G$ acts on $P$ via an affine representation, then it is a subtle question to determine whether there exists a $G$-invariant regular lattice triangulation of $P$. As another example, in Ehrhart theory, one associates to $P$ an invariant $h^{*}(P;t)$, which is a polynomial with nonnegative integer coefficients. If a group $G$ acts on $P$ via an affine representation $\rho$, then there exists a natural invariant $h^{*}(P,\rho;t)$ whose coefficients are virtual representations of $G$ and which equals $h^{*}(P;t)$ when the group action is trivial. It is a subtle question to determine whether the coefficients of $h^{*}(P,\rho;t)$ have nonnegative coefficients when expressed as sums of the classes of irreducible representations. The goal of this paper is to study and relate several such problems.
Ehrhart theory is the study of the enumeration of lattice points in lattice polytopes. Equivariant Ehrhart theory is a generalization of Ehrhart theory that takes into account the action of a finite group acting via affine transformations on the underlying lattice and preserving the polytope. It was introduced in [Sta11a] and studied in [ASVM20, ASVM21, EKS23, CHK23]. Before stating our results, we first recall some background on equivariant Ehrhart theory. Let $M\cong\mathbb{Z}^{d}$ be a lattice of rank $d$ and write $M_{k}:=M\otimes_{\mathbb{Z}}k$ for a field $k$. Let $P\subset M_{\mathbb{R}}$ be a $d$-dimensional lattice polytope. Let $\operatorname{Aff}(M)$ be the group of affine transformations of $M$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation of $G$ (see Section 2.2 and Section 2.3)). We assume that $P$ is $G$-invariant. Let $R(G)$ be the complex representation ring of $G$ (see Section 2.3). Then $R(G)$ is the free $\mathbb{Z}$-module with basis given by the classes $[V]$ in $R(G)$, as $V$ varies over all irreducible finite-dimensional $\mathbb{C}G$-modules, and an element of $R(G)$ with nonnegative integer coefficients with respect to this basis is called effective. We say that a power series in $R(G)[[t]]$ is effective if all its coefficients are effective.
@@
[…]
@@
(see [Sta11a, Theorem 7.7]) Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Consider the following conditions:
@@
Then the following implications hold: (1) $\Rightarrow$ (2) $\Rightarrow$ (3).
@@
We now present Theorem 1.4, a generalization of Theorem 1.1 that is the main result of the paper. Let $C_{P}\subset\widetilde{M}_{\mathbb{R}}$ be the cone generated by $P\times\{1\}$, and consider the graded semigroup algebra $S_{P}:=\mathbb{C}[C_{P}\cap\widetilde{M}$], with grading induced by $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$. For any positive integer $m$, there is a natural bijection between elements of $C_{P}\cap\operatorname{pr}^{-1}(m)$ and elements of $P\cap\frac{1}{m}M$, taking $(u,m)\mapsto\frac{1}{m}u$. In particular, $G$ acts on $S_{P}$ as a graded $\mathbb{C}$-algebra and $\operatorname{Ehr}(P,\rho;t)$ is the equivariant Hilbert series of $S_{P}$ (see Section 4.1). By a result of Hochster [Hoc72], $S_{P}$ is a Cohen-Macaulay ring. A linear system of parameters (l.s.o.p.) is a collection of $d+1$ homogeneous elements $F_{1},\ldots,F_{d+1}$ of degree $1$ in $S_{P}$ such that $S_{P}/(F_{1},\ldots,F_{d+1})$ is a finite dimensional $\mathbb{C}$-vector space (see Definition 4.3 and Theorem 4.4). If $F_{1},\ldots,F_{d+1}$ is a l.s.o.p. for $S_{P}$, then the ideal $(F_{1},\ldots,F_{d+1})$ is $G$-invariant if and only if the $\mathbb{C}$-vector space $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}$ is $G$-invariant if and only if we may consider $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]$ as an element of $R(G)$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $\widetilde{M}=M\oplus\mathbb{Z}$ and let $C_{P}$ be the cone over $P\times\{1\}$ in $\widetilde{M}_{\mathbb{R}}$ with corresponding graded semigroup algebra $S_{P}=\mathbb{C}[C_{P}\cap\widetilde{M}$]. Consider the following conditions:
@@
    There exists a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{P}$ such that $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}$] in $R(G)$.
@@
Then the following implications hold: (1) $\Rightarrow$ (2) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (5). Moreover, if there exists a $G$-invariant (not necessarily regular) lattice triangulation of $P$, then (4) holds.
When $G$ acts trivially, all the conditions of Theorem 1.4 are well-known to hold. See, for example, [DLRS10, Lemma 2.3.15] for condition (1) and [Sta76, Sta80] for conditions (3) and (4). We observe that condition (1) in Theorem 1.4 is often easier to verify in practice than conditions (2) or (3). In Proposition 3.12 we give an explicit criterion to guarantee that (1) holds i.e. a $G$-invariant regular lattice triangulation of $P$ exists. Using the notion of a translative action studied in [DR18], [DD21], [BD24] and [DD23] (see Definition 3.13) we deduce two corollaries for refining and gluing invariant lattice polyhedral subdivisions to construct invariant lattice triangulations (see Corollary 3.17 and Corollary 3.18 respectively). For example, we deduce that (1) holds when $G=\mathbb{Z}/2\mathbb{Z}$ acts on a centrally symmetric lattice polytope (see Example 3.21). On the other hand, we can view condition (5) in Theorem 1.4 as an obstruction that can be used to construct examples where conditions (1)-(3) fail.
For $G$-invariant lattice triangulations with some special additional structure, combinatorial formulas for the coefficients of $h^{*}(P,\rho;t)$ as explicit permutation representations were given in [EKS23, Theorem 3.30] (see, also, [EKS23, Theorem 3.32]). When the action of $G$ is translative and there exists a $G$-invariant unimodular lattice triangulation of $P$, D’alì and Delucchi showed in independent work that the equivariant Ehrhart series can be expressed as a rational function with denominator $(1-t)^{d+1}$ and gave a criterion which guarantees that the corresponding numerator is effective [DD23, Theorem 5.2], with applications to order polytopes and Lipschitz polytopes of posets ([DD23, Theorem 6.7] and [DD23, Corollary 7.12] respectively).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $N$ be a positive integer divisible by $|G|$. Then there exists a $G$-invariant regular triangulation of $P$ with vertices in $\frac{1}{N}M$.
@@
On the one hand, the definition is motivated by the non-equivariant case when $h^{*}_{N}(P,\rho;t)$ has an interpretation as the weighted Ehrhart polynomial of the $N$th dilate of the pyramid over $P$ (see Lemma 2.33 and [Sta08]). On the other hand, the definition is also motivated by the Ehrhart theory of rational polytopes where denominators with factors of the form $1-t^{N}$ (rather than 1t1-t) naturally appear (see Remark 2.2). In fact, we leave it as an open problem to extend the theory of this paper to the case when $P$ is a rational (rather than a lattice) polytope (cf. [CHK23]). We recover the invariant $h^{*}(P,\rho_{N};t)$ from $h_{N}^{*}(P,\rho;t)$ by extracting all terms with integer (rather than just rational) exponents in $t$ (see Lemma 2.33).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $N$ be a positive integer. Consider the following conditions:
@@
Then the following implications hold: (1) $\Rightarrow$ (2) $\Rightarrow$ (3).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $Q\subset P$ be a $G$-invariant (not necessarily full dimensional) lattice polytope. Suppose there exists a $G$-invariant lattice triangulation of $P$ that restricts to a $G$-invariant lattice triangulation of $Q$. Then $h^{*}(Q,\rho_{Q};t)\leq h^{*}(P,\rho;t)$.
@@
[…]
Secondly, in the statement of condition (2) we consider a polynomial $f\in\mathbb{C}[M]$ that is $G$-invariant. This implies that the hypersurface $\{f=0\}\subset\operatorname{Spec}\mathbb{C}[M]$ is $G$-invariant. However, the converse is false. More precisely, there may be a nontrivial character $\chi:G\to\mathbb{C}$ such that $g\cdot f=\chi(g)f$ for all $g$ in $G$, and then $\{f=0\}$ is $G$-invariant but $f$ is not. In particular, references to $G$-invariant hypersurfaces in [Sta11a] and [Sta11b] should be references to $G$-invariant polynomials. We leave open the problem of extending the theory to deal with all $G$-invariant hypersurfaces.
@@ 2.1. Ehrhart theory of rational polytopes
for some $0\leq\lambda_{k}<1$, $m_{i},m_{i}^{\prime}\in\mathbb{Z}$ and $\beta\in\mathbb{R}$. Note that $\lambda_{k}$ is determined by $\lambda_{k}=\beta a_{i}=-\beta a_{i}\in\mathbb{R}/\mathbb{Z}$ for any $i\in I_{k}$. Since $\gcd(a_{1},\ldots,a_{r})=1$, the condition $2\beta a_{i}=0\in\mathbb{R}/\mathbb{Z}$ implies that either $\beta=0\in\mathbb{R}/\mathbb{Z}$ or $\beta=\frac{1}{2}\in\mathbb{R}/\mathbb{Z}$. If $\beta=0\in\mathbb{R}/\mathbb{Z}$, then $\lambda_{k}=0$ for all $k$ and $\bar{u}=0$. Assume that $\beta=\frac{1}{2}\in\mathbb{R}/\mathbb{Z}$. Then $\lambda_{k}=\frac{1}{2}$ if $a_{i}$ is odd for some/any $i\in I_{k}$, and $\lambda_{k}=0$ otherwise. The only solution is $u=\frac{1}{2}\sum_{i:a_{i}\textrm{ odd }}(e_{i}+f_{i})$.